### v1

# Algorithms Trained on Normal Chest X-rays Can Predict Health Insurance Types

Abstract
Artificial intelligence is revealing what medicine never intended to encode. Deep vision models, trained on chest X-rays, can now detect not only disease but also invisible traces of social inequality. In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.67 on MIMIC-CXR-JPG, 0.68 on CheXpert). The signal persists even when age, race, and sex are controlled for, and remains detectable when the model is trained exclusively on a single racial group. Patch-based occlusion reveals that the signal is diffuse rather than localized, embedded in the upper and mid-thoracic regions. This suggests that deep networks may be internalizing subtle traces of clinical environments, equipment differences, or care pathways; learning socioeconomic segregation itself. These findings challenge the assumption that medical images are neutral biological data. By uncovering how models perceive and exploit these hidden social signatures, this work reframes fairness in medical AI: the goal is no longer only to balance datasets or adjust thresholds, but to interrogate and disentangle the social fingerprints embedded in clinical data itself.
keywords:
spurious correlation, health insurance, medical images
Data and Code Availability
The code is published on code.
Institutional Review Board (IRB)
IRB approval is not required in our study
1 Introduction
Artificial intelligence (AI) has become increasingly part of our everyday life. However, experts acknowledge concerns regarding AI fairness, specifically the disparate outcomes of an AI model across different societal subpopulations chexclusion. Such considerations are paramount given that the outcomes of AI-facilitated decision-making are directly correlated with an individual’s well-being. Recent studies shortcut; LnL; gender have demonstrated that deep learning methods are prone to exploit spurious correlation from shortcut features, which are predictive features that do not correspond to disease-specific patterns shortcuts_cause_bias. The data distribution may also be polluted by multiple political, historical, or socioeconomic factors that exist in human society, containing unethical information that could be realized by AI algorithms and used by them. For example, dissecting showed that a widely used commercial health prediction system treated different racial groups unequally since black patients were likely to spend less on healthcare than white patients, which might be correlated to hindered access to healthcare facilities for black patients. Therefore, researchers have been investigating the equality of AI algorithms treating people with different demographic data attributes (such as sex, race, age) across various modalities. gender_imbalance; noLLMfree.
While shortcut features could be obvious and easy to identify, such as chest tube or watermarks on a medical image; we are more concerned with those that are imperceptible and unidentifiable to humans, such as demographic information in a chest X-ray. Demographic information could even introduce significant societal bias if they act as a shortcut feature. A leading study gichoya2022ai showed that the race of patients is strongly detectable from chest X-rays and other studies localized race features on medical images to highlight how AI detect race from chest X-rays konate2025interpretability; salvado2024localisation. adleberg2022predicting discussed the hidden health insurance information in the chest X-ray images. However, the study included patients of all ages in MIMIC-IV, overlooking the fact that people older than 65 are automatically enrolled in public insurance. Moreover, they did not exclude the chest X-rays with thoracic diseases, which is another confounding factor of health insurance status.
In this finding paper, we demonstrated the potential of AI algorithms to detect health insurance types from chest X-ray images. The lower performance of AI model in Chest X-ray disease diagnosis chexclusion for low income patients with public insurance and the underdiagnosis bias against those patients has been documented in seyyed2021underdiagnosis; chexclusion; bahre2025underdiagnosis, necessitating further investigation to validate the presence of health insurance information within chest X-ray images. Therefore, our contribution is as below:
  1. 1.
    To the best of our knowledge, this is the first in-depth analysis that reveals that deep vision models (DenseNet densenet, SwinTransformer swin, MedMamba medmamba) can predict health insurance types from normal chest X-ray images (without thoracic diseases) in the MIMIC-CXR-JPG and CheXpert dataset. This suggests that chest X-rays (CXRs) contain latent socioeconomic information about patients that AI models can potentially leverage.
  2. 2.
    We localized health insurance type information in chest X-ray images and demonstrated that there is likely more information in the upper two-third part of a chest X-ray.
  3. 3.
    From close analysis between health insurance types and other demographic features (e.g. age, race, sex), we found that these common demographic features do not act as an essential mediator in our insurance type prediction task, which indicates that the deep vision models directly detect insurance type signals from chest X-ray images.
2 Methods
2.1 Datasets
In our experiment, two medical datasets were utilized: MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 and CheXpert chexpert. MIMIC-CXR-JPG is an extended image dataset for MIMIC-IV v3.0 mimic1; mimic2, including 377,110 chest X-ray images in total. We utilized MIMIC-CXR-JPG and linked it to MIMIC-IV v3.0 data by subject ID to attain patients’ health insurance type. There are six health insurance categories in MIMIC-IV v3.0: Medicaid, Medicare, Private, Self-pay, No charge, and Other. On the other hand, CheXpert includes 224,316 chest X-rays from 65,240 patients from centers of Stanford Hospital. The recent CheXpert Plus paper chambon2024chexpert provides additional health insurance information for each patient, in which there are three categories: Medicaid, Medicare, and Private. As for CheXpert, we utilized the JPG-downsized version due to computational limits. (See appendix A for detailed description.) For both datasets, we chose patients with Medicaid, Medicare, and Private insurance labels for analysis, further merging Medicaid and Medicare into the ”Public” label, in order to mainly inspect the disparity between people receiving public and private insurance. If a patient is covered by more than one type of health insurance in their lifetime, we chose the health insurance type that inferred the lowest socioeconomic status for the patient (i.e. Medicaid << Medicare << Public). As for the preprocessing of other demographic attributes, the age attribute was separated into three groups: 1-39, 40-49, 50-64. The division points were established to ensure an equitable distribution of patient counts across each group. As for race, we selected the two major ethnic groups in the datasets: White and Black, and the remaining ethnicities were pooled as ”Others”.
To decrease confounding factors that possibly affect our analysis, we only included 41,571 and 10,220 chest X-ray images in MIMIC-CXR-JPG and CheXpert, respectively, that matched the following three conditions: (1) Patient’s age under 65. (2) Image labeled as “no finding” and without ”support devices”. (3) Frontal view image (including posteroanterior/anteroposterior views). In the first condition, we excluded patients that were automatically covered by the public insurance, and focused our analyses on economically underprivileged patients in the public insurance group; in the second condition, we excluded patients with thoracic diseases to focus on how normal chest X-ray itself conveyed socioeconomic information; in the last condition, only frontal view images were considered to simplify our analyses. We randomly divided our extracted dataset into 0.8/0.1/0.1 splits for training, validation, and testing set, respectively, and ensured that patients in each split set did not overlap.
2.2 Models and Metrics
In our study, we utilized three different vision models: DenseNet121 densenet, SwinTransformer swin, MedMamba medmamba. Each of them represents a specific type of deep vision model. As for further training details, please refer to appendix B. As for health insurance type classifying performance, AUC (area under receiver operating characteristic curve) was utilized as our main metric.
3 Experiments
3.1 Health Insurance Prediction from CXRs
Setting In our first experiment, we have trained and tested our three chosen model: DenseNet121, SwinTransformer, and MedMamba on both MIMIC-CXR-JPG and CheXpert, respectively, to predict the health insurance type of each patient.
Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with MedMamba having the best average AUC (0.6669) on the insurance predicting task and SwinTransformer performed the worst (0.6261). On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6761. Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in SwinTransformer and MedMamba. The result suggested that the models are not simply guessing and that chest X-rays contain certain insurance information for the model to learn from.
[…]
tab:mainAUC DenseNet Swin MedMamba MIMIC 0.6385 0.6261 0.6669 CheXpert 0.6761 0.6168 0.6188
Table 1: AUC Performance on Insurance Prediction
3.2 Localization of insurance information on Xray - Patch-based training
Setting We investigated whether insurance information could be localized to a particular anatomical region or image patch with two different approaches. We divided each chest X-ray into 3x3 squared cells of approximately equal size. (Most grids are in size of 150x150, as some grids are sized 148x150 or 148x148 due to indivisible width and height.) We trained a health insurance classification model based on MedMamba using two different approaches. (1) Remove-One-Patch: Select one of the nine patches, and remove all information from that patch by setting pixel values within the patch to zero. (2) Keep-One-Patch: Select one of the nine patches, and remove all information from the image by setting pixel values to zero except for the pixels within the patch. Visualization of the two approaches are shown in Figure 1. MIMIC-CXR-JPG was chosen as the main dataset for this experiment due to its larger data size.
Results Figure 1 reveals the AUC performance on the insurance type prediction task for the Remove-One-Patch and Keep-One-Patch experiement. The lowest performance (0.6332) occurs when we remove the center part of the image, while the highest performance (0.6572) appears when pixels at the mid-lower corner are set to zero. Conversely, the highest performance (0.6541) in Keep-one-patch experiment lands in the top-left corner when it is the only part that is preserved in an image. AUC performance drops to the lowest (0.6063) when only the bottom-right part of the image is retained. While the pulmonary area is located at the middle of a chest radiograph, these results suggest that the health insurance information is not only inferred from lungs. Combining the two experiments, the upper two-third part seems to contain the most health insurance type information, while the bottom part of a chest radiograph contains the least of it.
[…]
fig:patch_based_1 Refer to caption
Figure 1: (a) The left image provides an example of the Remove-One-Patch method. The right image shows insurance type prediction AUC when only the corresponding patch area in the 3x3 grid is removed. (b) The left image provides an example of the Keep-One-Patch method. The right image shows insurance type prediction AUC when only the corresponding patch area in the 3x3 grid is retained.
3.3 Experiments on Demographic Mediators
Setting Based on the premise that deep learning models can effectively discern health insurance types from chest X-ray images, a pertinent inquiry arises: Is the insurance classification directly inferred, or is it mediated through demographic features like age, sex, or race? To answer this, we conducted an insurance detection experiment using other demographics. Specifically, we trained tree-based and clustering machine learning models including Random Forest randomforest, Decision Tree decisiontree, XGBoost xgboost, CatBoost catboost, and KNN knn to predict health insurance type from a combination of demographic attributes: age, sex, and race. The hyperparameter is tuned by GridSearchCV from the Scikit-Learn package. In addition, given the observation that the portion of Black people covered by private insurance is disproportionately lower than that in the White, we also conducted an experiment in which we trained, validated, and tested a health insurance type classifier only on patients that are reported as white. In this part of experiment, MIMIC-CXR-JPG and MedMamba were also chosen as the main dataset and deep vision model.
Result Table 2 shows the performance of multiple machine learning models to predict health insurance from the combination of age, race, and sex attributes. Among the five machine learning models (Random Forest, XGBoost, CatBoost, Decision Tree, KNN) that we have evaluated, XGBoost have the highest AUC value, but could only reach 0.5905. Therefore, we concluded that the results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6. Table 3 demonstrates the health insurance type prediction performance of the DenseNet121 model trained solely from White people in our dataset. We observed that the performance is relatively consistent with training our model with the entire dataset, except a noticeable drop DenseNet. In this experiment, we could still observe the superiority of MedMamba over the other two models. Our results indicate that the ability of deep vision models to predict health insurance types is not mediated by race features.
[…]
tab:MLforInsurance Model AUC Random Forest 0.5527 XGBoost 0.5905 CatBoost 0.5847 KNN 0.5475 Decision Tree 0.5527
Table 2: Health insurance type prediction performance across multiple machine learning methods given the combination of age, race, and sex attributes.
[…]
tab:race_stratified_AUC DenseNet Swin MedMamba 0.6047 0.6290 0.6679
Table 3: AUC of DenseNet121 trained on isolated White people in MIMIC-CXR-JPG
4 Discussion
Despite our results in AI algorithms predicting insurance information from chest X-ray images, there are several limitations in our work. First, we only chose one of the insurance types per person in their lifetime, which ignored the possible socioeconomic transition of a patient. Second, we did not explicitly show that health insurance information biases disease prediction in deep vision training, but since humans could easily learn the association between socioeconomic status and disease distribution, we should assume it being a simple association for a model to learn from. Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in DenseNet performance in isolated White people in MIMIC-CXR-JPG.
Our work serves as a calling to re-examine the current AI models trained on chest X-ray datasets as they may utilize socioeconomic information as shortcut feature for disease diagnosis. Moreover, we are looking forward to actively developing methods that ensure algorithms do not use these shortcut features. Although many features correlated with socioeconomic status may also be legitimate biological markers of a disease, we are aiming for techniques that can disentangle these signals and completely remove the influence of features that only predict socioeconomic status, while downgrading the contribution of features that are associated with both socioeconomic status and the disease of interest. In this way, we move the field toward developing truly robust and equitable clinical tools, rather than simply creating more powerful versions of biased systems.
5 Conclusion
We demonstrated that the state-of-the-art deep vision models learn health insurance type information from chest X-ray images, and that it is not mediated by common demographic factors. Despite the convenience of AI technology, we should utilize these deep learning models more cautiously when applying them to real-world medical tasks.
Appendix A Dataset description
MIMIC-IV v3.0 mimic1; mimic2 is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
On the other hand, CheXpert chexpert includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017. The recent CheXpert Plus paper chambon2024chexpert provides additional demographic information of each patient, including their health insurance type, race, sex, and age.
Appendix B Model architecture
DenseNet121 densenet comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model. SwinTransformer swin is a variant of vision transformer model, where self-attention is performed on shifted image windows. This idea largely decreases the memory complexity and reduces the downsides of separated image areas. In our study, the version of SwinV2-B* was utilized specifically. Lastly, MedMamba medmamba combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch. As for the classification, a single linear layer was added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
All of the experiments utilized Lion lion as their optimizer, and the learning rate was 1e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba. The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance. The batch size was 32 across our experiments, and we did not implement dropout or weight decay for regularization for simplicity. In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG. All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation. A cross-entropy loss was utilized for health insurance type classification tasks. We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset. The final performance was tested on the testing dataset with the chosen weights. All experiments were trained on a single NVIDIA GeForce RTX 3090 or RTX 6000 Ada GPU.

### v3

@@
 Deep vision models, trained on chest X-rays, can now detect not only disease but also invisible traces of social inequality.
-In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.67 on MIMIC-CXR-JPG, 0.68 on CheXpert).
+In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.70 on MIMIC-CXR-JPG, 0.68 on CheXpert).
 The signal persists even when age, race, and sex are controlled for, and remains detectable when the model is trained exclusively on a single racial group.
@@
 Setting In our first experiment, we have trained and tested our three chosen model: DenseNet121, SwinTransformer, and MedMamba on both MIMIC-CXR-JPG and CheXpert, respectively, to predict the health insurance type of each patient.
-Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with MedMamba having the best average AUC (0.6669) on the insurance predicting task and SwinTransformer performed the worst (0.6261).
-On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6761.
-Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in SwinTransformer and MedMamba.
+Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with DenseNet having the best average AUC (0.7007) on the insurance predicting task and SwinTransformer performed the worst (0.6182).
+On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6802.
+Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in all models.
 The result suggested that the models are not simply guessing and that chest X-rays contain certain insurance information for the model to learn from.
 […]
-tab:mainAUC DenseNet Swin MedMamba MIMIC 0.6385 0.6261 0.6669 CheXpert 0.6761 0.6168 0.6188
+tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6802 0.6063 0.6183
 Table 1: AUC Performance on Insurance Prediction
@@
 Results Figure 1 reveals the AUC performance on the insurance type prediction task for the Remove-One-Patch and Keep-One-Patch experiement.
-The lowest performance (0.6332) occurs when we remove the center part of the image, while the highest performance (0.6572) appears when pixels at the mid-lower corner are set to zero.
+The lowest performance (0.6386) occurs when we remove the center part of the image, while the highest performance (0.6624) appears when pixels at the bottom-right corner are set to zero.
 Conversely, the highest performance (0.6576) in Keep-one-patch experiment lands in the top-left corner when it is the only part that is preserved in an image.
-AUC performance drops to the lowest (0.6063) when only the bottom-right part of the image is retained.
+AUC performance drops to the lowest (0.6093) when only the mid-bottom part of the image is retained.
 While the pulmonary area is located at the middle of a chest radiograph, these results suggest that the health insurance information is not only inferred from lungs.
@@
 Therefore, we concluded that the results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6.
-Table 3 demonstrates the health insurance type prediction performance of the DenseNet121 model trained solely from White people in our dataset.
-We observed that the performance is relatively consistent with training our model with the entire dataset, except a noticeable drop DenseNet.
-In this experiment, we could still observe the superiority of MedMamba over the other two models.
+Table 3 demonstrates the health insurance type prediction performance of the MedMamba model trained solely from White people in our dataset.
+We observed that the performance is relatively consistent with training our model with the entire dataset.
+In this experiment, we could still observe the superiority of DenseNet over the other two models.
 Our results indicate that the ability of deep vision models to predict health insurance types is not mediated by race features.
@@
 […]
-tab:race_stratified_AUC DenseNet Swin MedMamba 0.6047 0.6290 0.6679
+tab:race_stratified_AUC DenseNet Swin MedMamba 0.6954 0.6154 0.6714
 Table 3: AUC of DenseNet121 trained on isolated White people in MIMIC-CXR-JPG
@@
 As for the classification, a single linear layer was added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
-All of the experiments utilized Lion lion as their optimizer, and the learning rate was 1e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba.
+All of the experiments utilized Lion lion as their optimizer, and the learning rate was 5e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba.
 The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance.

### v4

@@
 In this study, we show that state-of-the-art architectures (DenseNet121, SwinV2-B*, MedMamba) can predict a patient’s health insurance type, a strong proxy for socioeconomic status, from normal chest X-rays with significant accuracy (AUC $\approx$ 0.70 on MIMIC-CXR-JPG, 0.68 on CheXpert).
-The signal persists even when age, race, and sex are controlled for, and remains detectable when the model is trained exclusively on a single racial group.
+The signal was unlikely contributed by demographic features by our machine learning study combining age, race, and sex labels to predict health insurance types.
+The signal also remains detectable when the model is trained exclusively on a single racial group.
 Patch-based occlusion reveals that the signal is diffuse rather than localized, embedded in the upper and mid-thoracic regions.
@@
 By uncovering how models perceive and exploit these hidden social signatures, this work reframes fairness in medical AI: the goal is no longer only to balance datasets or adjust thresholds, but to interrogate and disentangle the social fingerprints embedded in clinical data itself.
+The codes are available at Link.
 keywords:
@@
 Artificial intelligence (AI) has become increasingly part of our everyday life.
-However, experts acknowledge concerns regarding AI fairness, specifically the disparate outcomes of an AI model across different societal subpopulations chexclusion.
+However, experts acknowledge concerns regarding AI fairness, specifically the disparate outcomes of an AI model across different societal subpopulations Seyyed-Kalantari et al.
+(2020).
 Such considerations are paramount given that the outcomes of AI-facilitated decision-making are directly correlated with an individual’s well-being.
-Recent studies shortcut; LnL; gender have demonstrated that deep learning methods are prone to exploit spurious correlation from shortcut features, which are predictive features that do not correspond to disease-specific patterns shortcuts_cause_bias.
+Recent studies Geirhos et al.
+(2020); Kim et al.
+(2019); Buolamwini and Gebru (2018) have demonstrated that deep learning methods are prone to exploit spurious correlation from shortcut features, which are predictive features that do not correspond to disease-specific patterns Banerjee et al.
+(2023).
 The data distribution may also be polluted by multiple political, historical, or socioeconomic factors that exist in human society, containing unethical information that could be realized by AI algorithms and used by them.
-For example, dissecting showed that a widely used commercial health prediction system treated different racial groups unequally since black patients were likely to spend less on healthcare than white patients, which might be correlated to hindered access to healthcare facilities for black patients.
+For example, Obermeyer et al.
+(2019) showed that a widely used commercial health prediction system treated different racial groups unequally since black patients were likely to spend less on healthcare than white patients, which might be correlated to hindered access to healthcare facilities for black patients.
 Therefore, researchers have been investigating the equality of AI algorithms treating people with different demographic data attributes (such as sex, race, age) across various modalities.
-gender_imbalance; noLLMfree.
+Larrazabal et al.
+(2020); Kumar et al.
+(2025).
 While shortcut features could be obvious and easy to identify, such as chest tube or watermarks on a medical image; we are more concerned with those that are imperceptible and unidentifiable to humans, such as demographic information in a chest X-ray.
 Demographic information could even introduce significant societal bias if they act as a shortcut feature.
-A leading study gichoya2022ai showed that the race of patients is strongly detectable from chest X-rays and other studies localized race features on medical images to highlight how AI detect race from chest X-rays konate2025interpretability; salvado2024localisation.
-adleberg2022predicting discussed the hidden health insurance information in the chest X-ray images.
-However, the study included patients of all ages in MIMIC-IV, overlooking the fact that people older than 65 are automatically enrolled in public insurance.
+A leading study Gichoya et al.
+(2022) showed that the race of patients is strongly detectable from chest X-rays and other studies localized race features on medical images to highlight how AI detect race from chest X-rays Konate et al.
+(2025); Salvado et al.
+(2024).
+Adleberg et al.
+(2022) discussed the hidden health insurance information in the chest X-ray images.
+However, the study included patients of all ages in MIMIC-IV, overlooking the fact that people older than 65 years old are automatically enrolled in public insurance.
 Moreover, they did not exclude the chest X-rays with thoracic diseases, which is another confounding factor of health insurance status.
 In this finding paper, we demonstrated the potential of AI algorithms to detect health insurance types from chest X-ray images.
-The lower performance of AI model in Chest X-ray disease diagnosis chexclusion for low income patients with public insurance and the underdiagnosis bias against those patients has been documented in seyyed2021underdiagnosis; chexclusion; bahre2025underdiagnosis, necessitating further investigation to validate the presence of health insurance information within chest X-ray images.
+The lower performance of AI model in Chest X-ray disease diagnosis Seyyed-Kalantari et al.
+(2020) for low income patients with public insurance and the underdiagnosis bias against those patients has been documented in Seyyed-Kalantari et al.
+(2021, 2020); Bahre et al.
+(2025), necessitating further investigation to validate the presence of health insurance information within chest X-ray images.
 Therefore, our contribution is as below:
   1. 1.
-    To the best of our knowledge, this is the first in-depth analysis that reveals that deep vision models (DenseNet densenet, SwinTransformer swin, MedMamba medmamba) can predict health insurance types from normal chest X-ray images (without thoracic diseases) in the MIMIC-CXR-JPG and CheXpert dataset.
+    To the best of our knowledge, this is the first in-depth analysis that reveals that deep vision models (DenseNet121 Huang et al.
+(2017), SwinV2-B* Liu et al.
+(2022), MedMamba Yue and Li (2024)) can predict health insurance types from normal chest X-ray images (without thoracic diseases) in the MIMIC-CXR-JPG and CheXpert dataset.
 This suggests that chest X-rays (CXRs) contain latent socioeconomic information about patients that AI models can potentially leverage.
   2. 2.
-    We localized health insurance type information in chest X-ray images and demonstrated that there is likely more information in the upper two-third part of a chest X-ray.
+    We demonstrated that these information distributed diffusely rather than localized and there is likely more information in the upper two-third part of a chest X-ray.
   3. 3.
@@
 2.1 Datasets
-In our experiment, two medical datasets were utilized: MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 and CheXpert chexpert.
-MIMIC-CXR-JPG is an extended image dataset for MIMIC-IV v3.0 mimic1; mimic2, including 377,110 chest X-ray images in total.
+In our experiment, two medical datasets were utilized: MIMIC-CXR-JPG Johnson et al.
+(2019a, b) and CheXpert Irvin et al.
+(2019).
+MIMIC-CXR-JPG is an extended image dataset for MIMIC-IV v3.0 Johnson et al.
+(2024, 2023), including 377,110 chest X-ray images in total.
 We utilized MIMIC-CXR-JPG and linked it to MIMIC-IV v3.0 data by subject ID to attain patients’ health insurance type.
@@
 On the other hand, CheXpert includes 224,316 chest X-rays from 65,240 patients from centers of Stanford Hospital.
-The recent CheXpert Plus paper chambon2024chexpert provides additional health insurance information for each patient, in which there are three categories: Medicaid, Medicare, and Private.
+The recent CheXpert Plus paper Chambon et al.
+(2024) provides additional health insurance information for each patient, in which there are three categories: Medicaid, Medicare, and Private.
 As for CheXpert, we utilized the JPG-downsized version due to computational limits.
-(See appendix A for detailed description.) For both datasets, we chose patients with Medicaid, Medicare, and Private insurance labels for analysis, further merging Medicaid and Medicare into the ”Public” label, in order to mainly inspect the disparity between people receiving public and private insurance.
-If a patient is covered by more than one type of health insurance in their lifetime, we chose the health insurance type that inferred the lowest socioeconomic status for the patient (i.e.
+(See appendix A for detailed dataset descriptions.) For both datasets, we chose patients with Medicaid, Medicare, and Private insurance labels for analysis, further merging Medicaid and Medicare into the ”Public” label, in order to mainly inspect the disparity between people receiving public and private insurance.
+If a patient is covered by more than one type of health insurance in their lifetime (in less than 8% of patients in MIMIC-IV), we chose the health insurance type that inferred the lowest socioeconomic status for the patient (i.e.
 Medicaid << Medicare << Public).
 As for the preprocessing of other demographic attributes, the age attribute was separated into three groups: 1-39, 40-49, 50-64.
-The division points were established to ensure an equitable distribution of patient counts across each group.
+The division points were established to ensure a relatively equitable distribution of patient counts across each group.
 As for race, we selected the two major ethnic groups in the datasets: White and Black, and the remaining ethnicities were pooled as ”Others”.
-To decrease confounding factors that possibly affect our analysis, we only included 41,571 and 10,220 chest X-ray images in MIMIC-CXR-JPG and CheXpert, respectively, that matched the following three conditions: (1) Patient’s age under 65.
+To decrease confounding factors that possibly affect our analysis, we only included 36,255 and 6,261 chest X-ray images in MIMIC-CXR-JPG and CheXpert, respectively, that matched the following three conditions: (1) Patient’s age under 65.
 (2) Image labeled as “no finding” and without ”support devices”.
@@
 We randomly divided our extracted dataset into 0.8/0.1/0.1 splits for training, validation, and testing set, respectively, and ensured that patients in each split set did not overlap.
+The detailed dataset demographic distribution is shown in 1
+Table 1: Dataset Demographics
+Train Valid Test
+(N=29,188) (N=3,439) (N=3,628)
+MIMIC-CXR-JPG
+Gender
+Male 14,241 (39.0%)* 1,675 (39.0%) 1,840 (35.0%)
+Female 14,947 (36.8%) 1,764 (36.3%) 1,788 (36.3%)
+Age Group
+1-39 7,926 (42.7%) 854 (47.0%) 949 (43.1%)
+40-49 6,907 (37.2%) 870 (38.7%) 858 (36.8%)
+50-64 14,355 (35.5%) 1,715 (32.5%) 1,821 (31.4%)
+Race
+White 17,132 (43.5%) 2,059 (40.1%) 2,052 (42.4%)
+Black 7,220 (27.1%) 808 (33.7%) 935 (22.6%)
+Others 4,836 (33.9%) 572 (34.6%) 641 (33.5%)
+CheXpert
+Train Valid Test
+(N=5,007) (N=639) (N=615)
+Gender
+Male 2,892 (55.7%)* 396 (58.1%) 325 (56.0%)
+Female 2,115 (47.5%) 242 (46.3%) 290 (46.6%)
+Age Group
+1-39 1,886 (56.9%) 219 (64.4%) 229 (53.3%)
+40-49 1,086 (59.1%) 161 (57.8%) 154 (59.1%)
+50-64 2,035 (44.2%) 259 (42.1%) 232 (44.8%)
+Race
+White 2,519 (57.8%) 322 (58.4%) 298 (59.7%)
+Black 398 (24.6%) 59 (25.4%) 53 (22.6%)
+Others 2,090 (50.8%) 258 (54.3%) 264 (48.1%)
+* Proportion of patients with Private Insurance within that subgroup.
 2.2 Models and Metrics
-In our study, we utilized three different vision models: DenseNet121 densenet, SwinTransformer swin, MedMamba medmamba.
+In our study, we utilized three different vision models: DenseNet121 Huang et al.
+(2017), SwinV2-B* Liu et al.
+(2022), MedMamba Yue and Li (2024).
 Each of them represents a specific type of deep vision model.
-As for further training details, please refer to appendix B.
-As for health insurance type classifying performance, AUC (area under receiver operating characteristic curve) was utilized as our main metric.
+DenseNet121 Huang et al.
+(2017) comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model.
+SwinTransformer Liu et al.
+(2022) is a variant of vision transformer model, where self-attention is performed on shifted image windows.
+This idea largely decreases the memory complexity and reduces the downsides of separated image areas.
+In our study, the version of SwinV2-B* was utilized specifically.
+Lastly, MedMamba Yue and Li (2024) combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch.
+As for the classification, two linear layers were added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
+All of the experiments utilized Lion Liang et al.
+(2024) as their optimizer, and the learning rate was 5e-6 for DenseNet121, 1e-5 for SwinV2-B* and MedMamba if not otherwise mentioned.
+The batch size was 32 across MIMIC experiments, and 128, 128, 32, respectively for DenseNet121, SwinV2-B*, MedMamba in CheXpert experiments after hyperparameter search.
+We did not implement dropout or weight decay for regularization for simplicity.
+In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG.
+All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation in the training and validation process, while only Normalization was performed on testing images.
+A cross-entropy loss was utilized for health insurance type classification tasks.
+We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset, which is our main evaluation metric for insurance type prediction performance.
+The final performance was tested on the testing dataset with the chosen weights.
+All experiments were trained on a single NVIDIA GeForce RTX 3090, or H200 clusters.
 3 Experiments
 3.1 Health Insurance Prediction from CXRs
-Setting In our first experiment, we have trained and tested our three chosen model: DenseNet121, SwinTransformer, and MedMamba on both MIMIC-CXR-JPG and CheXpert, respectively, to predict the health insurance type of each patient.
-Results As Table 1 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance around 0.65, with DenseNet having the best average AUC (0.7007) on the insurance predicting task and SwinTransformer performed the worst (0.6182).
-On the other hand, the models trained by on CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6802.
-Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in all models.
-The result suggested that the models are not simply guessing and that chest X-rays contain certain insurance information for the model to learn from.
+Setting In our first experiment, we have trained and tested our three chosen model: DenseNet121, SwinV2-B*, and MedMamba on both MIMIC-CXR-JPG and CheXpert, respectively, to predict the health insurance type of each patient.
+We did an additional control experiment by random label perturbation (i.e.
+assigning label randomly to each image) on MIMIC-CXR-JPG to construct MIMIC-Random dataset.
+Results As Table 2 indicates, in MIMIC-CXR-JPG, all trained health insurance predicting models could achieve an AUC performance over 0.61, with DenseNet having the best average AUC (0.7007) on the insurance predicting task and SwinV2-B* performed the worst (0.6182).
+On the other hand, the models trained by CheXpert also had a better-than-random performance on the health insurance prediction, as the best performance is 0.6834.
+Notice that CheXpert contains less and smaller sized images compared to MIMIC-CXR-JPG, which may be the major cause of the AUC decrease in all three models.
+When the model was trained with the MIMIC-Random dataset, the AUC performance degraded to the level of random guess.
+The results supports our hypothesis that the models are not simply guessing health insurance types and that chest X-rays contain certain insurance information for the models to learn from.
 […]
-tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6802 0.6063 0.6183
+tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6834 0.6063 0.6183 MIMIC-Rd 0.5011 0.4811 0.5058
 Table 2: AUC Performance on Insurance Prediction
@@
 Visualization of the two approaches are shown in Figure 1.
-MIMIC-CXR-JPG was chosen as the main dataset for this experiment due to its larger data size.
+MIMIC-CXR-JPG was chosen as the main dataset for this experiment due to its larger dataset size.
 Results Figure 1 reveals the AUC performance on the insurance type prediction task for the Remove-One-Patch and Keep-One-Patch experiement.
-The lowest performance (0.6386) occurs when we remove the center part of the image, while the highest performance (0.6624) appears when pixels at the bottom-right corner are set to zero.
-Conversely, the highest performance (0.6576) in Keep-one-patch experiment lands in the top-left corner when it is the only part that is preserved in an image.
-AUC performance drops to the lowest (0.6093) when only the mid-bottom part of the image is retained.
-While the pulmonary area is located at the middle of a chest radiograph, these results suggest that the health insurance information is not only inferred from lungs.
-Combining the two experiments, the upper two-third part seems to contain the most health insurance type information, while the bottom part of a chest radiograph contains the least of it.
+The lowest performance (0.6508) occurs when we remove the right-upper part of the image, while the highest performance (0.6674) appears when pixels at the mid-upper patch are set to zero.
+Notably there is no localized pattern in this experiment that locates the insurance type information, and removing each patch only hurts very little performance.
+Conversely, there is a relatively clear localized pattern in Keep-one-patch experiment.
+The highest performance (0.6378) lands in the top-left corner when it is the only preserved part in an image.
+AUC performance drops to the lowest (0.5916) when only the bottom-right part of the image is retained.
+Combining the results, it is suggested that the health insurance type information is diffusely distributed in the chest X-ray image (by the Remove-One-Patch experiment), but relatively more information concentrated at the upper two-third part on the image (by the Keep-One-Patch experiement).
 […]
@@
 To answer this, we conducted an insurance detection experiment using other demographics.
-Specifically, we trained tree-based and clustering machine learning models including Random Forest randomforest, Decision Tree decisiontree, XGBoost xgboost, CatBoost catboost, and KNN knn to predict health insurance type from a combination of demographic attributes: age, sex, and race.
+Specifically, we trained tree-based and clustering machine learning models including Random Forest Breiman (2001), Decision Tree Quinlan (1986), XGBoost Chen and Guestrin (2016), CatBoost Prokhorenkova et al.
+(2018), and KNN Fix (1985) to predict health insurance type from a combination of demographic attributes: age, sex, and race.
 The hyperparameter is tuned by GridSearchCV from the Scikit-Learn package.
-In addition, given the observation that the portion of Black people covered by private insurance is disproportionately lower than that in the White, we also conducted an experiment in which we trained, validated, and tested a health insurance type classifier only on patients that are reported as white.
-In this part of experiment, MIMIC-CXR-JPG and MedMamba were also chosen as the main dataset and deep vision model.
+In addition, given the disproportionately low private insurance coverage for Black patients compared to White patients in the dataset (1), we trained a health insurance classifier only on the White subgroup in MIMIC-CXR-JPG and CheXpert datasets.
+If the health insurance type prediction is mediated by the racial features, we should observe a performance degradation on the experiments.
+The learning rate for Lion optimizer was adjusted to 1e-4 in Swin-V2-B* trained with isolated White people CheXpert.
 Result Table 3 shows the performance of multiple machine learning models to predict health insurance from the combination of age, race, and sex attributes.
 Among the five machine learning models (Random Forest, XGBoost, CatBoost, Decision Tree, KNN) that we have evaluated, XGBoost have the highest AUC value, but could only reach 0.5905.
-Therefore, we concluded that the results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6.
-Table 3 demonstrates the health insurance type prediction performance of the MedMamba model trained solely from White people in our dataset.
-We observed that the performance is relatively consistent with training our model with the entire dataset.
-In this experiment, we could still observe the superiority of DenseNet over the other two models.
-Our results indicate that the ability of deep vision models to predict health insurance types is not mediated by race features.
+These results do not indicate a strong correlation between health insurance types and the three demographic features, as all the AUCs are no higher than 0.6.
+Table 4 demonstrates the health insurance type prediction performance trained solely with White people in our datasets.
+We observed that models trained with only White people images from MIMIC-CXR-JPG still managed to learn the health insurance information and suffers from little AUC degradation.
+However, the performance dropped substantially when training solely with White people in CheXpert.
+Notice that there is only 2508 White people in the CheXpert training dataset, compared to 17,132 White people in the MIMIC-CXR-JPG training dataset.
+This huge population difference provides a possible explanation for the performance degradation.
+On the other hand, Densenet could still maintain a high AUC performance around 0.65, suggesting that racial features should not be a significant mediator in learning health insurance type.
+Our results indicate that the ability of deep vision models to predict health insurance types is not likely mediated by race features, at least not a significant contributor.
 […]
@@
 […]
-tab:race_stratified_AUC DenseNet Swin MedMamba 0.6954 0.6154 0.6714
-Table 3: AUC of DenseNet121 trained on isolated White people in MIMIC-CXR-JPG
+tab:race_stratified_AUC DenseNet Swin MedMamba MIMIC 0.6954 0.6154 0.6714 CheXpert 0.6535 0.5847 0.5808
+Table 4: AUC of DenseNet121 trained on isolated White people in MIMIC-CXR-JPG and CheXpert
 4 Discussion
+Interpreting these results requires considering what kinds of features AI models can extract from medical images.
+A useful conceptual framing distinguishes three types of features that medical-AI models can exploit: (1) biological features, directly reflecting anatomy or pathology; (2) social features, deriving from social determinants of health such as socioeconomic status or insurance types; and (3) hybrid features that intertwine both domains; for instance, age or smoking status, which combine biological and social dimensions.
+The ethical and technical goal is to minimize reliance on purely social features that reflect structural disadvantage rather than biology.
+Simply removing explicit demographic variables is insufficient as models often reconstruct them indirectly from other data.
+Fairness in medical AI therefore depends not only on data curation but on recognizing and attenuating hidden social signals embedded in the pixels themselves.
+Our study design makes this challenge unusually visible.
+We trained and evaluated models solely on radiographs labeled “no finding,” images that lacked any CheXpert/MIMIC-defined abnormalities (cardiomegaly, enlarged cardiomediastinum, pulmonary edema, consolidation, pneumonia, pleural effusion, pneumothorax, atelectasis, fracture, lung lesion or opacity, and pleural or device-related findings), ensuring that all images were radiographically normal, frontal in view, and from patients under 65 years old.
+This design precludes shortcut learning from obvious pathology: the model could not rely on disease-related cues that might correlate with insurance status, yet its performance remained well above chance.
+That persistence implies the model captured either minute physiologic correlates of social conditions or contextual artifacts introduced during image acquisition.
+Some of these contextual artifacts may arise from system-level differences (such as hospital site, imaging hardware, or acquisition protocols) that correlate with insurance coverage, meaning the model could be detecting institutional rather than biological patterns.
+However, our experiments showed the strongest results using the single-site MIMIC dataset, indicating that other factors contribute to this phenomenon.
+The remaining signal likely originates in patient-level physiology itself, shaped over time by social environment, nutrition, stress, and comorbidity.
+To examine where such subtle information might reside, we turned to spatial localization maps.
+Predictive regions were concentrated in the upper and mid-thorax, with minimal contribution from the lung bases.
+This pattern is unlikely to arise from artifacts such as tubes, leads, or devices, and may instead reflect subtle physiologic correlates.
+The upper and mid-thorax encompass the heart, great vessels, ribs, spine, and soft tissues; structures that may encode subtle, population-level differences linked to socioeconomic context.
+Variations in bone density, soft-tissue distribution, or vascular morphology shaped by chronic stress, nutrition, and healthcare access could all leave faint but detectable traces.
+The model’s focus on these regions aligns with prior work suggesting that AI systems can infer age, sex, and even mortality risk from texture and posture cues invisible to the human eye Ieki et al.
+(2022); Raghu et al.
+(2021); Li et al.
+(2022).
+Nevertheless, imaging-acquisition factors (e.g., variations in positioning, exposure, or calibration that correlate with insurance type) remain plausible contributors.
+Disentangling these biological and contextual pathways will require controlled studies linking quantitative anatomic features and imaging parameters to socioeconomic indicators.
+If these cues are truly physiologic or contextual rather than demographic, they should persist across subgroups.
+Indeed, in experiments restricted to White participants, performance remained nearly unchanged.
+In another experiment predicting health insurance solely from combined demographic features also revealed low performance.
+These results suggest that demographic features is not acting as a direct proxy for insurance type in Chest X-rays.
+The model detects something more subtle; latent, physiologically mediated imprints of social conditions that escape coarse demographic categorization.
+This finding strengthens our argument that “invisible social fingerprints” can persist even when explicit social variables are absent, challenging the assumption that models trained on ostensibly neutral data are free from social bias.
 Despite our results in AI algorithms predicting insurance information from chest X-ray images, there are several limitations in our work.
-First, we only chose one of the insurance types per person in their lifetime, which ignored the possible socioeconomic transition of a patient.
+First, both the MIMIC-IV cohort we used and the current CheXpert-Plus dataset (as of February 2024) simplify a patient’s history by reporting only one insurance type.
+In MIMIC-IV, we specifically selected the insurance type associated with the lowest socioeconomic status.
+Both approaches fail to account for possible socioeconomic transitions during a patient’s lifetime.
+However, an analysis of the original MIMIC-IV dataset revealed that only around 8% of patients ever changed their insurance status, suggesting a limited impact by this transition factor.
 Second, we did not explicitly show that health insurance information biases disease prediction in deep vision training, but since humans could easily learn the association between socioeconomic status and disease distribution, we should assume it being a simple association for a model to learn from.
-Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in DenseNet performance in isolated White people in MIMIC-CXR-JPG.
+Based on past literature discovering social biases in medical image predictions Seyyed-Kalantari et al.
+(2021), this is a reasonable assumption.
+Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in SwinV2-B* and MedMamba performance in isolated White people in CheXpert.
 Our work serves as a calling to re-examine the current AI models trained on chest X-ray datasets as they may utilize socioeconomic information as shortcut feature for disease diagnosis.
@@
 5 Conclusion
-We demonstrated that the state-of-the-art deep vision models learn health insurance type information from chest X-ray images, and that it is not mediated by common demographic factors.
+We demonstrated that the state-of-the-art deep vision models learn health insurance type information from chest X-ray images, and that it is not essentially mediated by common demographic factors.
 Despite the convenience of AI technology, we should utilize these deep learning models more cautiously when applying them to real-world medical tasks.
 Appendix A Dataset description
-MIMIC-IV v3.0 mimic1; mimic2 is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
-On the other hand, CheXpert chexpert includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017.
-The recent CheXpert Plus paper chambon2024chexpert provides additional demographic information of each patient, including their health insurance type, race, sex, and age.
-Appendix B Model architecture
-DenseNet121 densenet comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model.
-SwinTransformer swin is a variant of vision transformer model, where self-attention is performed on shifted image windows.
-This idea largely decreases the memory complexity and reduces the downsides of separated image areas.
-In our study, the version of SwinV2-B* was utilized specifically.
-Lastly, MedMamba medmamba combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch.
-As for the classification, a single linear layer was added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
-All of the experiments utilized Lion lion as their optimizer, and the learning rate was 5e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba.
-The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance.
-The batch size was 32 across our experiments, and we did not implement dropout or weight decay for regularization for simplicity.
-In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG.
-All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation.
-A cross-entropy loss was utilized for health insurance type classification tasks.
-We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset.
-The final performance was tested on the testing dataset with the chosen weights.
-All experiments were trained on a single NVIDIA GeForce RTX 3090 or RTX 6000 Ada GPU.
+MIMIC-IV v3.0 Johnson et al.
+(2024, 2023) is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG Johnson et al.
+(2019a, b) is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
+On the other hand, CheXpert Irvin et al.
+(2019) includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017.
+The images were downsized to 390 x 320 in the downsized version.
+The recent CheXpert Plus paper Chambon et al.
+(2024) provides additional demographic information of each patient, including their health insurance type, race, sex, and age.

### v5

@@
 Artificial intelligence (AI) has become increasingly part of our everyday life.
-However, experts acknowledge concerns regarding AI fairness, specifically the disparate outcomes of an AI model across different societal subpopulations Seyyed-Kalantari et al.
-(2020).
+However, experts acknowledge concerns regarding AI fairness, specifically the disparate outcomes of an AI model across different societal subpopulations chexclusion.
 Such considerations are paramount given that the outcomes of AI-facilitated decision-making are directly correlated with an individual’s well-being.
-Recent studies Geirhos et al.
-(2020); Kim et al.
-(2019); Buolamwini and Gebru (2018) have demonstrated that deep learning methods are prone to exploit spurious correlation from shortcut features, which are predictive features that do not correspond to disease-specific patterns Banerjee et al.
-(2023).
+Recent studies shortcut; LnL; gender have demonstrated that deep learning methods are prone to exploit spurious correlation from shortcut features, which are predictive features that do not correspond to disease-specific patterns shortcuts_cause_bias.
 The data distribution may also be polluted by multiple political, historical, or socioeconomic factors that exist in human society, containing unethical information that could be realized by AI algorithms and used by them.
-For example, Obermeyer et al.
-(2019) showed that a widely used commercial health prediction system treated different racial groups unequally since black patients were likely to spend less on healthcare than white patients, which might be correlated to hindered access to healthcare facilities for black patients.
+For example, dissecting showed that a widely used commercial health prediction system treated different racial groups unequally since black patients were likely to spend less on healthcare than white patients, which might be correlated to hindered access to healthcare facilities for black patients.
 Therefore, researchers have been investigating the equality of AI algorithms treating people with different demographic data attributes (such as sex, race, age) across various modalities.
-Larrazabal et al.
-(2020); Kumar et al.
-(2025).
+gender_imbalance; noLLMfree.
 While shortcut features could be obvious and easy to identify, such as chest tube or watermarks on a medical image; we are more concerned with those that are imperceptible and unidentifiable to humans, such as demographic information in a chest X-ray.
 Demographic information could even introduce significant societal bias if they act as a shortcut feature.
-A leading study Gichoya et al.
-(2022) showed that the race of patients is strongly detectable from chest X-rays and other studies localized race features on medical images to highlight how AI detect race from chest X-rays Konate et al.
-(2025); Salvado et al.
-(2024).
-Adleberg et al.
-(2022) discussed the hidden health insurance information in the chest X-ray images.
+A leading study gichoya2022ai showed that the race of patients is strongly detectable from chest X-rays and other studies localized race features on medical images to highlight how AI detect race from chest X-rays konate2025interpretability; salvado2024localisation.
+adleberg2022predicting discussed the hidden health insurance information in the chest X-ray images.
 However, the study included patients of all ages in MIMIC-IV, overlooking the fact that people older than 65 years old are automatically enrolled in public insurance.
@@
 In this finding paper, we demonstrated the potential of AI algorithms to detect health insurance types from chest X-ray images.
-The lower performance of AI model in Chest X-ray disease diagnosis Seyyed-Kalantari et al.
-(2020) for low income patients with public insurance and the underdiagnosis bias against those patients has been documented in Seyyed-Kalantari et al.
-(2021, 2020); Bahre et al.
-(2025), necessitating further investigation to validate the presence of health insurance information within chest X-ray images.
+The lower performance of AI model in Chest X-ray disease diagnosis chexclusion for low income patients with public insurance and the underdiagnosis bias against those patients has been documented in seyyed2021underdiagnosis; chexclusion; bahre2025underdiagnosis, necessitating further investigation to validate the presence of health insurance information within chest X-ray images.
 Therefore, our contribution is as below:
   1. 1.
-    To the best of our knowledge, this is the first in-depth analysis that reveals that deep vision models (DenseNet121 Huang et al.
-(2017), SwinV2-B* Liu et al.
-(2022), MedMamba Yue and Li (2024)) can predict health insurance types from normal chest X-ray images (without thoracic diseases) in the MIMIC-CXR-JPG and CheXpert dataset.
+    To the best of our knowledge, this is the first in-depth analysis that reveals that deep vision models (DenseNet121 densenet, SwinV2-B* swin, MedMamba medmamba) can predict health insurance types from normal chest X-ray images (without thoracic diseases) in the MIMIC-CXR-JPG and CheXpert dataset.
 This suggests that chest X-rays (CXRs) contain latent socioeconomic information about patients that AI models can potentially leverage.
@@
 2.1 Datasets
-In our experiment, two medical datasets were utilized: MIMIC-CXR-JPG Johnson et al.
-(2019a, b) and CheXpert Irvin et al.
-(2019).
-MIMIC-CXR-JPG is an extended image dataset for MIMIC-IV v3.0 Johnson et al.
-(2024, 2023), including 377,110 chest X-ray images in total.
+In our experiment, two medical datasets were utilized: MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 and CheXpert chexpert.
+MIMIC-CXR-JPG is an extended image dataset for MIMIC-IV v3.0 mimic1; mimic2, including 377,110 chest X-ray images in total.
 We utilized MIMIC-CXR-JPG and linked it to MIMIC-IV v3.0 data by subject ID to attain patients’ health insurance type.
@@
 On the other hand, CheXpert includes 224,316 chest X-rays from 65,240 patients from centers of Stanford Hospital.
-The recent CheXpert Plus paper Chambon et al.
-(2024) provides additional health insurance information for each patient, in which there are three categories: Medicaid, Medicare, and Private.
+The recent CheXpert Plus paper chambon2024chexpert provides additional health insurance information for each patient, in which there are three categories: Medicaid, Medicare, and Private.
 As for CheXpert, we utilized the JPG-downsized version due to computational limits.
@@
 2.2 Models and Metrics
-In our study, we utilized three different vision models: DenseNet121 Huang et al.
-(2017), SwinV2-B* Liu et al.
-(2022), MedMamba Yue and Li (2024).
-Each of them represents a specific type of deep vision model.
-DenseNet121 Huang et al.
-(2017) comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model.
-SwinTransformer Liu et al.
-(2022) is a variant of vision transformer model, where self-attention is performed on shifted image windows.
+In our study, we utilized three different vision models: DenseNet121 densenet, SwinV2-B* swin, MedMamba medmamba. Each of them represents a specific type of deep vision model.
+DenseNet121 densenet comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model.
+SwinTransformer swin is a variant of vision transformer model, where self-attention is performed on shifted image windows.
 This idea largely decreases the memory complexity and reduces the downsides of separated image areas.
 In our study, the version of SwinV2-B* was utilized specifically.
-Lastly, MedMamba Yue and Li (2024) combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch.
+Lastly, MedMamba medmamba combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch.
 As for the classification, two linear layers were added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
-All of the experiments utilized Lion Liang et al.
-(2024) as their optimizer, and the learning rate was 5e-6 for DenseNet121, 1e-5 for SwinV2-B* and MedMamba if not otherwise mentioned.
+All of the experiments utilized Lion lion as their optimizer, and the learning rate was 5e-6 for DenseNet121, 1e-5 for SwinV2-B* and MedMamba if not otherwise mentioned.
 The batch size was 32 across MIMIC experiments, and 128, 128, 32, respectively for DenseNet121, SwinV2-B*, MedMamba in CheXpert experiments after hyperparameter search.
@@
 […]
-tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6834 0.6063 0.6183 MIMIC-Rd 0.5011 0.4811 0.5058
+tab:mainAUC DenseNet Swin MedMamba MIMIC 0.7007 0.6182 0.6684 CheXpert 0.6834 0.6063 0.6183 MIMIC-Rd 0.5058 0.5011 0.4811
 Table 2: AUC Performance on Insurance Prediction
@@
 To answer this, we conducted an insurance detection experiment using other demographics.
-Specifically, we trained tree-based and clustering machine learning models including Random Forest Breiman (2001), Decision Tree Quinlan (1986), XGBoost Chen and Guestrin (2016), CatBoost Prokhorenkova et al.
-(2018), and KNN Fix (1985) to predict health insurance type from a combination of demographic attributes: age, sex, and race.
+Specifically, we trained tree-based and clustering machine learning models including Random Forest randomforest, Decision Tree decisiontree, XGBoost xgboost, CatBoost catboost, and KNN knn to predict health insurance type from a combination of demographic attributes: age, sex, and race.
 The hyperparameter is tuned by GridSearchCV from the Scikit-Learn package.
@@
 Variations in bone density, soft-tissue distribution, or vascular morphology shaped by chronic stress, nutrition, and healthcare access could all leave faint but detectable traces.
-The model’s focus on these regions aligns with prior work suggesting that AI systems can infer age, sex, and even mortality risk from texture and posture cues invisible to the human eye Ieki et al.
-(2022); Raghu et al.
-(2021); Li et al.
-(2022).
+The model’s focus on these regions aligns with prior work suggesting that AI systems can infer age, sex, and even mortality risk from texture and posture cues invisible to the human eye ieki2022deep; raghu2021deep; li2022deep.
 Nevertheless, imaging-acquisition factors (e.g., variations in positioning, exposure, or calibration that correlate with insurance type) remain plausible contributors.
@@
 Second, we did not explicitly show that health insurance information biases disease prediction in deep vision training, but since humans could easily learn the association between socioeconomic status and disease distribution, we should assume it being a simple association for a model to learn from.
-Based on past literature discovering social biases in medical image predictions Seyyed-Kalantari et al.
-(2021), this is a reasonable assumption.
+Based on past literature discovering social biases in medical image predictions seyyed2021underdiagnosis, this is a reasonable assumption.
 Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in SwinV2-B* and MedMamba performance in isolated White people in CheXpert.
@@
 Appendix A Dataset description
-MIMIC-IV v3.0 Johnson et al.
-(2024, 2023) is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG Johnson et al.
-(2019a, b) is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
-On the other hand, CheXpert Irvin et al.
-(2019) includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017.
+MIMIC-IV v3.0 mimic1; mimic2 is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
+On the other hand, CheXpert chexpert includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017.
 The images were downsized to 390 x 320 in the downsized version.
-The recent CheXpert Plus paper Chambon et al.
-(2024) provides additional demographic information of each patient, including their health insurance type, race, sex, and age.
+The recent CheXpert Plus paper chambon2024chexpert provides additional demographic information of each patient, including their health insurance type, race, sex, and age.